of every 8 hot dogs Lester sold 3 had mustard what percent did not have mustard?
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

3/8 had mustard, so 1-3/8 = 5/8 did not have mustard

5/8 converts to (5/8)*100 = 62.5%

So 62.5% of the hot dogs sold did not have mustard.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by stanbon(75887)

Ans: 5/8 = 62.5%
------------------
Cheers,
Stan H.